OpenStudy (henriquerb):

Please, help me with that =) Integrate dx/(a+bx^n)^((n+1)/n) with a different from 0} Thank you guys

OpenStudy (anonymous):

\(\displaystyle\int\frac{dx}{(a+bx^n)^{(n+1)/n}}\) with \(a\neq0\) ?

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

In the denominator, take \[\left(a+bx^n\right)^{(n+1)/n}=\left(x^n \right)^{(n+1)/n}\left(\frac{a}{x^n}+b\right)^{(n+1)/n}=x^{n+1}\left(\frac{a}{x^n}+b\right)^{(n+1)/n}\] Then substitute \(u=\dfrac{a}{x^n}+b\).

OpenStudy (anonymous):

\[\begin{align*}\int\frac{dx}{\left(a+bx^n\right)^{(n+1)/n}}&=\int\frac{dx}{x^{n+1}\left(\dfrac{a}{x^n}+b\right)^{(n+1)/n}}\\\\ &=-\frac{1}{an}\int\frac{du}{u^{(n+1)/n}} \end{align*}\]

OpenStudy (anonymous):

Notice that \(\dfrac{n+1}{n}\neq1\) for any \(n\), so you can immediately use the power rule and not have to worry about the antiderivative form of \(\dfrac{1}{u}\).

OpenStudy (henriquerb):

Thank you very much for your help.